OpenStudy (anonymous):

The line x + y = 2 intersects curve 2(x-1)^2 - (y+1)^2 =8 at the points A and B. Find the equation of the perpendicular bisector of the line joining A and B

OpenStudy (anonymous):

U have (-5,7) and (3,-1) -> slope 1 so ur bisector will have slope -1 Calculate mid point (x,y) and that plus the slope is your line.

OpenStudy (anonymous):

hmm equation of line is x + y =2 \(y = 2 - x \) for point intersecting points on curve we can substitute the y value in curve's equation \[2(x-1)^2 - (2-x+1)^2 = 8 \] \[2(x-1)^2 - (1 -x)^2 = 8 \] \[2(x^2 +1 -2x) - (1 + x^2 - 2x) = 8 \] \[x^2 + 1 - 2x - 8 = 0 \] Now you have a quadratic in x solve it you will get 2 points and similarly form a quadratic in y for getting y points the points that you will have will be same as estudier's for finding the the mid points \( x_{mid},y_{mid} = \frac{x_1 + x_2}{2},\frac{y_1+y_2}{2}\) Now mid point will be (-1 , 3)

OpenStudy (anonymous):

solve x by cross multiply right

OpenStudy (anonymous):

Now the line the perpendicular line should pass through (-1,3) And will have minus of reciprocal of slope of the line x + y =2 (slope of line x +y = 2 is -1 ) hence, slope of perpendicular line should be 1 y = x + c to determine c we will put the point that this line pass through we have (-1,3) as that point 3 = -1 +c c = 4 hence, we have an equation y = x +4

OpenStudy (anonymous):

lol i think cant you know

OpenStudy (anonymous):

You will have to solve the quadratic through the quadratic formula or factor by group

OpenStudy (anonymous):

in this case?

OpenStudy (anonymous):

but I am for if there is a better method available to determine point of intersection it's lengthy one that I did

OpenStudy (anonymous):

The quadratic equation we have is \[x^2 - 2x - 7 = 0 \]...hmm can anyone spot any errors in my solution I think I messed up somewhere